# Stochastic Cahn–Hilliard Equations from One-Dimensional Ising–Kac–Kawasaki Dynamics

### 31 Mar 2026 v1

@@ 4.1 One-block replacement
On each reference block $\Lambda_{\ell}$, we consider the bounded local observable $\Psi:\{-1,1\}^{\mathbb{Z}}\to\mathbb{R}$ with
$\mathrm{supp}(\Psi)\subset\Lambda_{r}:=\{-r,\dots,r\}$
for some $r\geq 1$. Using the local canonical Gibbs measure $\mu_{\gamma,\ell,m}^{\mathrm{can}}(d\eta)$, for every $\ell\geq r$, we define
$\Phi_{\Psi,\ell}^{\gamma}(m):=\int\Psi(\eta)\,\mu_{\gamma,\ell,m}^{\mathrm{can}}(d\eta),\quad m\in\mathcal{M}_{\ell}.$
We have the following one-block estimate.
@@
For every smooth test function $J\in C^{\infty}(\mathbb{T})$, as $\ell=\ell(\gamma)\to\infty,\gamma\ell(\gamma)\to 0$,
$\lim_{\ell_{\gamma}\to\infty}\limsup_{\gamma\ell_{\gamma}\to 0}\lim_{N\to\infty}\mathbb{E}_{f_{N,0}}\left|\int_{0}^{T}\frac{1}{N}\sum_{x\in\Lambda_{N}}J\!\left(\frac{x}{N}\right)\Bigl(\tau_{x}\Psi(\sigma(t))-\Phi_{\Psi,\ell}^{\gamma}(m_{x}^{\ell}(\sigma(t)))\Bigr)\,dt\right|=0.$
@@
Let $f_{N,t}$ be the density of the law of $\sigma(t)$ with respect to the canonical Gibbs measure $\mu_{N,\gamma,\beta}$. By reversibility and entropy dissipation,
$\frac{d}{dt}H_{N}(f_{N,t}\mid\mu_{N,\gamma,\beta})=-4N^{2}\sum_{x\in\Lambda_{N}}I_{x,x+1}^{N,\gamma}(f_{N,t}),$
where
$I_{x,x+1}^{N,\gamma}(f):=\frac{1}{4}\int c_{x,x+1}^{\gamma}(\sigma)\Bigl(\sqrt{f(\sigma^{x,x+1})}-\sqrt{f(\sigma)}\Bigr)^{2}\,\mu_{N,\gamma,\beta}(d\sigma).$
Hence
$H_{N}(f_{N,t}\mid\mu_{N,\gamma,\beta})\leq CN,\qquad\int_{0}^{T}\sum_{x\in\Lambda_{N}}I_{x,x+1}^{N,\gamma}(f_{N,t})\,dt\leq\frac{C}{N}.$
@@
$H_{N}(\bar{f}_{N}\mid\mu_{N,\gamma,\beta})\leq CN,\qquad\sum_{x\in\Lambda_{N}}I_{x,x+1}^{N,\gamma}(\bar{f}_{N})\leq\frac{C}{N}.$ (23)
@@
$Q_{N}^{\ell}(d\theta,d\eta):=d\theta\,\beta_{\theta}^{N,\ell}(d\eta)=\frac{1}{2N+1}\sum_{x\in\Lambda_{N}}\delta_{x/N}(d\theta)\,(\pi_{x,\ell})_{*}(\bar{f}_{N}\mu_{N,\gamma,\beta})(d\eta),\quad(\theta,\eta)\in\mathbb{T}\times\{-1,1\}^{\Lambda_{\ell}}.$
Passing to a subsequence if necessary, we may assume
@@
To prove this, write a global configuration as $\sigma=(\eta,\xi)$, where $\eta$ denotes the block variables and $\xi$ the exterior variables. Then $g_{x,\ell}(\eta)$ is obtained by conditional averaging of $\bar{f}_{N}(\eta,\xi)$ over $\xi$, and the exchange $\eta\mapsto\eta^{j,j+1}$ corresponds exactly to the global exchange on the bond $(x+j,x+j+1)$.
@@
$\sum_{|j|\leq\ell-1}\int_{\mathbb{T}}I_{j,j+1}^{(\ell),\gamma}\!\left(\frac{d\beta_{\theta}^{N,\ell}}{d\mu_{\gamma,\ell}^{\mathrm{gc}}}\right)\,d\theta\leq C_{\ell}\sum_{|j|\leq\ell-1}\frac{1}{N}\sum_{x\in\Lambda_{N}}I_{x+j,x+j+1}^{N,\gamma}(\bar{f}_{N}).$
@@
$\sum_{|j|\leq\ell-1}\frac{1}{N}\sum_{x\in\Lambda_{N}}I_{x+j,x+j+1}^{N,\gamma}(\bar{f}_{N})\leq C_{\ell}^{(4)}\sum_{b\in\Lambda_{N}}I_{b,b+1}^{N,\gamma}(\bar{f}_{N}).$
@@
$\sum_{|j|\leq\ell-1}\int_{\mathbb{T}}I_{j,j+1}^{(\ell),\gamma}\!\left(\frac{d\beta_{\theta}^{N,\ell}}{d\mu_{\gamma,\ell}^{\mathrm{gc}}}\right)\,d\theta\leq\frac{C_{\ell}}{N}\to 0.$
@@
$\int_{\sum_{\ell}}\Bigl(\Psi(\eta)-\Phi_{\Psi,\ell}^{\gamma}(m^{\ell}(\eta))\Bigr)\,\beta_{\theta}^{\ell}(d\eta)=0.$
@@
$\lim_{N\to\infty}\int J(\theta)\Bigl(\Psi(\eta)-\Phi_{\Psi,\ell}^{\gamma}(m^{\ell}(\eta))\Bigr)dQ_{N}^{\ell}=\int_{\mathbb{T}}J(\theta)\left[\int\Bigl(\Psi(\eta)-\Phi_{\Psi,\ell}^{\gamma}(m^{\ell}(\eta))\Bigr)\,\beta_{\theta}^{\ell}(d\eta)\right]d\theta=0.$
@@
$\lim_{N\to\infty}\mathbb{E}_{f_{N,0}}\left[\int_{0}^{T}\frac{1}{N}\sum_{x\in\Lambda_{N}}J\!\left(\frac{x}{N}\right)\Bigl(\tau_{x}\Psi(\sigma(t))-\Phi_{\Psi,\ell}^{\gamma}(m_{x}^{\ell}(\sigma(t)))\Bigr)\,dt\right]=0.$
Since $\Psi$ and $\Phi_{\Psi,\ell}^{\gamma}$ are uniformly bounded, the random variables above are uniformly bounded in $L^{\infty}$, so convergence of expectations implies convergence of the expectation of the absolute value. Finally, letting $\ell\to\infty$ with $\gamma\ell\to 0$ gives the claim. ∎
@@ 4.2 Two-block replacement for the conserved field
We first show the follow local $H^{-1}$ bound for bond differences.
@@
There exists a constant $C>0$, depending only on $c_{-}$ and $c_{+}$, such that for every $z\in\Lambda_{N}$ and every bounded function $f:\Sigma_{N}\to\mathbb{R}$,
$\bigl|\langle A_{z},f\rangle_{\mu_{N,\gamma,\beta}}\bigr|\leq C\,\sqrt{D_{z}^{N,\gamma}(f)}.$
Consequently, for every family of coefficients $(b_{z})_{z\in\Lambda_{N}}$, the function $W(\sigma):=\sum_{z\in\Lambda_{N}}b_{z}\,A_{z}(\sigma)$ satisfies
$\|W\|_{\mathcal{H}^{-1}_{N}}^{2}\leq\frac{C}{N^{2}}\sum_{z\in\Lambda_{N}}|b_{z}|^{2},$ (25)
where
$\|W\|_{\mathcal{H}^{-1}_{N}}^{2}:=\sup_{f}\Bigl\{2\langle W,f\rangle_{\mu_{N,\gamma,\beta}}-N^{2}D_{N}^{N,\gamma}(f)\Bigr\},\quad D_{N}^{N,\gamma}(f):=\sum_{z\in\Lambda_{N}}D_{z}^{N,\gamma}(f).$
@@
$(L_{z}^{N,\gamma}g)(\sigma):=c_{z}(\sigma)\bigl(g(\sigma^{z,z+1})-g(\sigma)\bigr)$
be the local exchange generator on that bond. We claim that there exists a bounded function $h_{z}:\Sigma_{N}\to\mathbb{R}$ such that $L_{z}^{N,\gamma}h_{z}=A_{z}.$ If $\sigma^{z,z+1}=\sigma$, then $\sigma(z)=\sigma(z+1)$, hence $A_{z}(\sigma)=\sigma(z+1)-\sigma(z)=0,$ and there is no constraint on $h_{z}$ on this orbit. If $\sigma^{z,z+1}\neq\sigma$, then the orbit consists of the pair $\{\sigma,\sigma^{z,z+1}\},$ and $A_{z}(\sigma^{z,z+1})=-A_{z}(\sigma).$ Choose arbitrarily one representative $\sigma$ from the orbit and set $h_{z}(\sigma)=0$. Then define
$h_{z}(\sigma^{z,z+1}):=\frac{A_{z}(\sigma)}{c_{z}(\sigma)}.$
By construction, we have
$(L_{z}^{N,\gamma}h_{z})(\sigma)=c_{z}(\sigma)\bigl(h_{z}(\sigma^{z,z+1})-h_{z}(\sigma)\bigr)=A_{z}(\sigma),$
For the swapped configuration $\sigma^{z,z+1}$, using detailed balance and the antisymmetry of $A_{z}$, the same identity holds automatically. Hence $L_{z}^{N,\gamma}h_{z}=A_{z}$ on the whole state space.
@@
$|h_{z}(\sigma^{z,z+1})-h_{z}(\sigma)|=\left|\frac{A_{z}(\sigma)}{c_{z}(\sigma)}\right|\leq\frac{2}{c_{-}}.$ (26)
@@
$\langle L_{z}^{N,\gamma}g,f\rangle_{\mu_{N,\gamma,\beta}}=-\frac{1}{2}\int c_{z}(\sigma)\,\bigl(g(\sigma^{z,z+1})-g(\sigma)\bigr)\bigl(f(\sigma^{z,z+1})-f(\sigma)\bigr)\,\mu_{N,\gamma,\beta}(d\sigma).$
@@
$\langle A_{z},f\rangle_{\mu_{N,\gamma,\beta}}=-\frac{1}{2}\int c_{z}(\sigma)\,\bigl(h_{z}(\sigma^{z,z+1})-h_{z}(\sigma)\bigr)\bigl(f(\sigma^{z,z+1})-f(\sigma)\bigr)\,\mu_{N,\gamma,\beta}(d\sigma).$
@@
$\leq$ $\left(\frac{1}{2}\int c_{z}(\sigma)\,\bigl(h_{z}(\sigma^{z,z+1})-h_{z}(\sigma)\bigr)^{2}\,d\mu_{N,\gamma,\beta}\right)^{1/2}\left(\frac{1}{2}\int c_{z}(\sigma)\,\bigl(f(\sigma^{z,z+1})-f(\sigma)\bigr)^{2}\,d\mu_{N,\gamma,\beta}\right)^{1/2}$
$\leq$ $C\sqrt{D_{z}^{N,\gamma}(f)},$
which proves the first claim. Then, by the local bound just proved, we have
$|\langle W,f\rangle_{\mu_{N,\gamma,\beta}}|\leq\sum_{z\in\Lambda_{N}}|b_{z}|\,|\langle A_{z},f\rangle_{\mu_{N,\gamma,\beta}}|\leq C\sum_{z\in\Lambda_{N}}|b_{z}|\,\sqrt{D_{z}^{N,\gamma}(f)}\leq N^{2}D_{N}^{N,\gamma}(f)+\frac{C}{N^{2}}\sum_{z\in\Lambda_{N}}|b_{z}|^{2}.$
Using the variational definition of the $\mathcal{H}^{-1}_{N}$-norm, we conclude (25). This proves the lemma. ∎
Theorem 6 (Two-block estimate).
For every smooth test function $J\in C^{\infty}(\mathbb{T})$,
$\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\left|\int_{0}^{T}\frac{1}{N}\sum_{x\in\Lambda_{N}}J\!\left(\frac{x}{N}\right)\Bigl(m_{x}^{\ell_{\gamma}}(\sigma(t))-m_{x}^{L_{\gamma}}(\sigma(t))\Bigr)\,dt\right|=0.$
@@
We split the argument into two steps.
Step 1: from $m_{x}^{\ell}$ to the sliding average of $\ell$-blocks. Define
$\widetilde{m}_{x}^{\ell,L}(\sigma):=\frac{1}{2L+1}\sum_{|y|\leq L}m_{x+y}^{\ell}(\sigma).$
Then
@@
Hence it suffices to prove
$\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\left|\int_{0}^{T}\frac{1}{N}\sum_{x\in\Lambda_{N}}J\!\left(\frac{x}{N}\right)\Bigl(m_{x}^{\ell}(\sigma(t))-\widetilde{m}_{x}^{\ell,L}(\sigma(t))\Bigr)\,dt\right|=0.$
Fix $y>0$. By definition,
$m_{x+y}^{\ell}-m_{x}^{\ell}=\frac{1}{2\ell+1}\sum_{u=-\ell}^{\ell}\bigl(\sigma_{x+y+u}-\sigma_{x+u}\bigr).$
Note that the exchange between $x+u$ and $x+y+u$ can be realized telescopically by the exchanges between neighboring sites from $x+u$ to $x+y+u$. Using the telescoping identity
$\sigma_{x+y+u}-\sigma_{x+u}=\sum_{k=0}^{y-1}\bigl(\sigma_{x+u+k+1}-\sigma_{x+u+k}\bigr),$
we obtain
$m_{x+y}^{\ell}-m_{x}^{\ell}=\frac{1}{2\ell+1}\sum_{u=-\ell}^{\ell}\sum_{k=0}^{y-1}\bigl(\sigma_{x+u+k+1}-\sigma_{x+u+k}\bigr).$
The same formula holds for $y<0$ after reversing the path.
Thus
$\frac{1}{N}\sum_{x\in\Lambda_{N}}J\!\left(\frac{x}{N}\right)\bigl(m_{x+y}^{\ell}-m_{x}^{\ell}\bigr)=\sum_{z\in\Lambda_{N}}b_{z}^{N,\ell,y}\bigl(\sigma(z+1)-\sigma(z)\bigr),$
for deterministic coefficients $b_{z}^{N,\ell,y}$. Averaging over $|y|\leq L$, we obtain
$\frac{1}{N}\sum_{x\in\Lambda_{N}}J\!\left(\frac{x}{N}\right)\bigl(m_{x}^{\ell}-\widetilde{m}_{x}^{\ell,L}\bigr)=\sum_{z\in\Lambda_{N}}b_{z}^{N,\ell,L}\bigl(\sigma(z+1)-\sigma(z)\bigr),$
with coefficients satisfying the counting bound $\sum_{z\in\Lambda_{N}}|b_{z}^{N,\ell,L}|^{2}\leq C_{J}\,\frac{L}{N}.$ Let
$W_{N,\ell,L}(\sigma):=\sum_{z\in\Lambda_{N}}b_{z}^{N,\ell,L}\bigl(\sigma_{z+1}-\sigma_{z}\bigr).$
Then Lemma 8 with $W=W_{N,\ell,L}$ yields that
$\mathbb{E}_{f_{N,0}}\left[\left(\int_{0}^{T}W_{N,\ell,L}(\sigma(t))\,dt\right)^{2}\right]\leq C\,T\,\|W_{N,\ell,L}\|_{\mathcal{H}^{-1}_{N}}^{2},$
Then by the variational characterization (25) of the $\mathcal{H}^{-1}$-norm and Young’s inequality,
$\|W_{N,\ell,L}\|_{\mathcal{H}^{-1}_{N}}^{2}\leq\frac{C}{N^{2}}\sum_{z\in\Lambda_{N}}|b_{z}^{N,\ell,L}|^{2}\leq C_{J}\,\frac{L}{N^{3}}.$
Hence
$\mathbb{E}_{f_{N,0}}\left[\left(\int_{0}^{T}W_{N,\ell,L}(\sigma(t))\,dt\right)^{2}\right]\leq C_{J}\,T\,\frac{L}{N^{3}}.$
In particular, since $L/N\to 0$,
$\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\left|\int_{0}^{T}\frac{1}{N}\sum_{x\in\Lambda_{N}}J\!\left(\frac{x}{N}\right)\Bigl(m_{x}^{\ell}(\sigma(t))-\widetilde{m}_{x}^{\ell,L}(\sigma(t))\Bigr)\,dt\right|=0.$
Step 2: from the sliding average to the larger block average. By elementary averaging,
$\widetilde{m}_{x}^{\ell,L}=\frac{1}{2L+1}\sum_{|y|\leq L}\frac{1}{2\ell+1}\sum_{|u|\leq\ell}\sigma_{x+y+u}=\sum_{v\in\mathbb{Z}}w_{\ell,L}(v)\sigma_{x+v},$
@@
$m_{x}^{L}=\sum_{|v|\leq L}\frac{1}{2L+1}\sigma_{x+v}.$
A direct comparison of the weights yields
$\sup_{\sigma\in\Sigma_{N}}\sup_{x\in\Lambda_{N}}|\widetilde{m}_{x}^{\ell,L}(\sigma)-m_{x}^{L}(\sigma)|\leq C\,\frac{\ell}{L}.$

### 05 Apr 2026 v2

@@ 4.1 One-block replacement
Let $\Psi:\{-1,1\}^{\mathbb{Z}}\to\mathbb{R}$ be a bounded local observable with $\operatorname{supp}(\Psi)\subset\Lambda_{r}$ for some $1\leq r<\ell_{\gamma}$, and let
$\Phi_{\Psi,\ell}^{\gamma}(m):=\int\Psi(\eta)\,\mu_{\gamma,\ell,m}^{\mathrm{can}}(d\eta),\qquad m\in\mathcal{M}_{\ell}.$
Under the block-scale assumptions above, for every bounded test function $J\in L^{\infty}(\mathbb{T})$,
$\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\left[\int_{0}^{T}\varepsilon_{\gamma}\sum_{i\in\Lambda_{N}}|J(\varepsilon_{\gamma}i)|\Bigl|\tau_{i}\Psi\bigl(\sigma^{N}(\alpha_{\gamma}^{-1}t)\bigr)-\Phi_{\Psi,\ell_{\gamma}}^{\gamma}\bigl(m_{i}^{\ell_{\gamma}}(\sigma^{N}(\alpha_{\gamma}^{-1}t))\bigr)\Bigr|\,dt\right]=0.$
@@
Let $f_{N,t}$ be the density of the law of $\sigma^{N}(t/\alpha_{\gamma})$ with respect to the canonical Gibbs measure $\mu_{N,\gamma,\beta}$. The entropy dissipation estimate in Lemma 8 yields that
$\frac{d}{dt}H_{N}(f_{N,t}\mid\mu_{N,\gamma,\beta})=-4\alpha_{\gamma}^{-1}I_{N}^{\gamma},$
Since $H_{N}(f_{N,t}\mid\mu_{N,\gamma,\beta})\leq CN$,
$4\alpha_{\gamma}^{-1}\int_{0}^{T}I_{N}^{\gamma}(f_{N,t})\,dt\leq H_{N}(f_{N,0}\mid\mu_{N,\gamma,\beta})\leq C_{H}N,$
The Scaling Assumption $(\mathbf{B})$ yields that $\alpha_{\gamma}N\to 0.$. Hence
$\int_{0}^{T}I_{N}^{\gamma}(f_{N,t})\,dt\leq C\alpha_{\gamma}N\to 0.$
@@
$H_{N}(\bar{f}_{N}\mid\mu_{N,\gamma,\beta})\leq CN,\qquad\sum_{x\in\Lambda_{N}}I_{x,x+1}^{N,\gamma}(\bar{f}_{N})\leq C\alpha_{\gamma}N.$ (26)
@@
$Q_{N}^{\ell}(d\theta,d\eta):=d\theta\,\beta_{\theta}^{N,\ell}(d\eta)=\varepsilon_{\gamma}\sum_{x\in\Lambda_{N}}\delta_{x/N}(d\theta)\,(\pi_{x,\ell})_{*}(\bar{f}_{N}\mu_{N,\gamma,\beta})(d\eta),\quad(\theta,\eta)\in\mathbb{T}\times\{-1,1\}^{\Lambda_{\ell}}.$
Passing to a subsequence if necessary, we assume
@@
Then $g_{x,\ell}(\eta)$ is obtained by conditional averaging of $\bar{f}_{N}(\eta,\xi)$ over $\xi$, and the exchange $\eta\mapsto\eta^{j,j+1}$ corresponds exactly to the global exchange on the bond $(x+j,x+j+1)$.
@@
$\sum_{|j|\leq\ell-1}\int_{\mathbb{T}}I_{j,j+1}^{(\ell),\gamma}\!\left(\frac{d\beta_{\theta}^{N,\ell}}{d\mu_{\gamma,\ell}^{\mathrm{gc}}}\right)\,d\theta\leq C_{\ell}\sum_{|j|\leq\ell-1}\varepsilon_{\gamma}\sum_{x\in\Lambda_{N}}I_{x+j,x+j+1}^{N,\gamma}(\bar{f}_{N}).$
@@
$\sum_{|j|\leq\ell-1}\varepsilon_{\gamma}\sum_{x\in\Lambda_{N}}I_{x+j,x+j+1}^{N,\gamma}(\bar{f}_{N})\leq C_{\ell}^{(4)}\sum_{b\in\Lambda_{N}}I_{b,b+1}^{N,\gamma}(\bar{f}_{N}).$
@@
$\sum_{|j|\leq\ell-1}\int_{\mathbb{T}}I_{j,j+1}^{(\ell),\gamma}\!\left(\frac{d\beta_{\theta}^{N,\ell}}{d\mu_{\gamma,\ell}^{\mathrm{gc}}}\right)\,d\theta\leq C_{\ell}\alpha_{\gamma}N\to 0.$
@@
$\int_{\sum_{\ell}}|J(\theta)|\,\bigl|\Psi(\eta)-\Phi_{\Psi,\ell}^{\gamma}(m)\bigr|\beta_{\theta}^{\ell}(d\eta)=0.$
@@
$\lim_{N\to\infty}\int_{\sum_{\ell}}|J(\theta)|\,\bigl|\Psi(\eta)-\Phi_{\Psi,\ell}^{\gamma}(m)\bigr|dQ_{N}^{\ell}\leq\int_{\mathbb{T}}|J(\theta)|\left|\int\Bigl(\Psi(\eta)-\Phi_{\Psi,\ell}^{\gamma}(m^{\ell}(\eta))\Bigr)\,\beta_{\theta}^{\ell}(d\eta)\right|d\theta=0.$
@@
$\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\left[\int_{0}^{T}\varepsilon_{\gamma}\sum_{i\in\Lambda_{N}}|J(\varepsilon_{\gamma}i)|\Bigl|\tau_{i}\Psi\bigl(\sigma^{N}(\alpha_{\gamma}^{-1}t)\bigr)-\Phi_{\Psi,\ell_{\gamma}}^{\gamma}\bigl(m_{i}^{\ell_{\gamma}}(\sigma^{N}(\alpha_{\gamma}^{-1}t))\bigr)\Bigr|\,dt\right]=0.$
This proves the theorem. ∎
@@ 4.2 Two-block replacement for the conserved field
We first show the follow local $\mathcal{H}_{N}^{-1}$ bound for the Kawasaki exchange of bond differences $A_{z}(\sigma):=\sigma(z+1)-\sigma(z)$.
@@
There exists a constant $C>0$, such that for every $z\in\Lambda_{N}$ and every bounded function $f:\Sigma_{N}\to\mathbb{R}$,
$\bigl|\langle A_{z},f\rangle_{\mu_{N,\gamma,\beta}}\bigr|=\bigl|\langle\sigma(z+1)-\sigma(z),f(\sigma)\rangle_{\mu_{N,\gamma,\beta}}\bigr|\leq C\,\sqrt{D_{z}^{N,\gamma}(f)}.$
@@
$(L_{z}^{N,\gamma}g)(\sigma):=c_{z}(\sigma)\bigl(g(\sigma_{z,z+1})-g(\sigma)\bigr)$
be the local exchange generator on that bond. We claim that there exists a bounded function $h_{z}:\Sigma_{N}\to\mathbb{R}$ such that $L_{z}^{N,\gamma}h_{z}=A_{z}.$ If $\sigma_{z,z+1}=\sigma$, then $\sigma(z)=\sigma(z+1)$, hence $A_{z}(\sigma)=\sigma(z+1)-\sigma(z)=0,$ and there is no constraint on $h_{z}$ on this orbit. If $\sigma_{z,z+1}\neq\sigma$, then the orbit consists of the pair $\{\sigma,\sigma_{z,z+1}\},$ and $A_{z}(\sigma_{z,z+1})=-A_{z}(\sigma).$ Choose arbitrarily one representative $\sigma$ from the orbit and set $h_{z}(\sigma)=0$. Then define
$h_{z}(\sigma_{z,z+1}):=\frac{A_{z}(\sigma)}{c_{z}(\sigma)}.$
By construction of $h_{z}$, we have
$(L_{z}^{N,\gamma}h_{z})(\sigma)=c_{z}(\sigma)\bigl(h_{z}(\sigma_{z,z+1})-h_{z}(\sigma)\bigr)=A_{z}(\sigma),$
For the swapped configuration $\sigma_{z,z+1}$, using detailed balance and the antisymmetry of $A_{z}$, the same identity holds automatically. Hence $L_{z}^{N,\gamma}h_{z}=A_{z}$ on the whole state space.
@@
$|h_{z}(\sigma_{z,z+1})-h_{z}(\sigma)|=\left|\frac{A_{z}(\sigma)}{c_{z}(\sigma)}\right|\leq\frac{2}{c_{-}}.$ (28)
@@
$\langle L_{z}^{N,\gamma}g,f\rangle_{\mu_{N,\gamma,\beta}}=-\frac{1}{2}\int c_{z}(\sigma)\,\bigl(g(\sigma_{z,z+1})-g(\sigma)\bigr)\bigl(f(\sigma_{z,z+1})-f(\sigma)\bigr)\,\mu_{N,\gamma,\beta}(d\sigma).$
@@
$\langle A_{z},f\rangle_{\mu_{N,\gamma,\beta}}=-\frac{1}{2}\int c_{z}(\sigma)\,\bigl(h_{z}(\sigma_{z,z+1})-h_{z}(\sigma)\bigr)\bigl(f(\sigma_{z,z+1})-f(\sigma)\bigr)\,\mu_{N,\gamma,\beta}(d\sigma).$
@@
$\leq$ $\left(\frac{1}{2}\int c_{z}(\sigma)\,\bigl(h_{z}(\sigma_{z,z+1})-h_{z}(\sigma)\bigr)^{2}\,d\mu_{N,\gamma,\beta}\right)^{1/2}\left(\frac{1}{2}\int c_{z}(\sigma)\,\bigl(f(\sigma_{z,z+1})-f(\sigma)\bigr)^{2}\,d\mu_{N,\gamma,\beta}\right)^{1/2}$
$\leq$ $C\sqrt{D_{z}^{N,\gamma}(f)}.$
This proves the lemma. ∎
Theorem 6 (Weak two-block estimate).
Under the block-scale assumptions above, for every bounded test function $J\in L^{\infty}(\mathbb{T})$,
$\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\left|\int_{0}^{T}\varepsilon_{\gamma}\sum_{i\in\Lambda_{N}}J(\varepsilon_{\gamma}i)\Bigl(m_{i}^{\ell_{\gamma}}\bigl(\sigma^{N}(\alpha_{\gamma}^{-1}t)\bigr)-m_{i}^{L_{\gamma}}\bigl(\sigma^{N}(\alpha_{\gamma}^{-1}t)\bigr)\Bigr)\,dt\right|=0.$ (29)
@@
We split the proof into two steps.
Step 1: Small block versus sliding average of small blocks. Define the sliding average of $\ell$-blocks
$\widetilde{m}_{x}^{\ell,L}(\sigma):=\frac{1}{2L+1}\sum_{|y|\leq L}m_{x+y}^{\ell}(\sigma),$
and set
$W_{\gamma,\ell,L}^{(1)}(\alpha_{\gamma}^{-1}t):=\varepsilon_{\gamma}\sum_{x\in\Lambda_{N}}J\left(\varepsilon_{\gamma}x\right)\Bigl(m_{x}^{\ell}(\alpha_{\gamma}^{-1}t)-\widetilde{m}_{x}^{\ell,L}(\alpha_{\gamma}^{-1}t)\Bigr).$
By definition of $\widetilde{m}_{x}^{\ell,L}$,
@@
For each fixed $y>0$,
$m_{x+y}^{\ell}-m_{x}^{\ell}=\frac{1}{2\ell+1}\sum_{u=-\ell}^{\ell}\bigl(\sigma(x+y+u)-\sigma(x+u)\bigr).$
Using the telescoping identity
$\sigma(x+y+u)-\sigma(x+u)=\sum_{k=0}^{y-1}\bigl(\sigma(x+u+k+1)-\sigma(x+u+k)\bigr).$
The same formula holds for $y<0$ after reversing the path. Hence every difference $m_{x+y}^{\ell}-m_{x}^{\ell}$ can be written as a finite sum of nearest-neighbour bond gradients.
By the telescoping representation, we obtains a bond decomposition
$W_{\gamma,\ell,L}^{(1)}(\sigma)=\sum_{z\in\Lambda_{N}}b_{z}^{\gamma,\ell,L}\bigl(\sigma(z+1)-\sigma(z)\bigr),$
where the coefficients $b_{z}^{\gamma,\ell,L}$ satisfy the counting estimate $\sum_{z\in\Lambda_{N}}|b_{z}^{\gamma,\ell,L}|^{2}\leq C\|J\|^{2}_{\infty}\frac{L}{N}.$
We now estimate the additive functional generated by $W_{\gamma,\ell,L}^{(1)}$. The local $\mathcal{H}_{N}^{-1}$ estimate in Lemma 11 yields that for bond differences,
$\Big|\langle\sigma(z+1)-\sigma(z),f(\sigma)\rangle_{\mu_{N,\gamma,\beta}}\Big|\leq C\,\sqrt{D_{z}^{N,\gamma}(f)}.$
Therefore
$\bigl|\langle W_{\gamma,\ell,L}^{(1)},f\rangle_{\mu_{N,\gamma,\beta}}\bigr|=\Big|\sum_{z\in\Lambda_{N}}b_{z}^{\gamma,\ell,L}\langle\sigma(z+1)-\sigma(z),f(\sigma)\rangle_{\mu_{N,\gamma,\beta}}\Big|\leq C\sum_{z\in\Lambda_{N}}|b_{z}^{\gamma,\ell,L}|\sqrt{D^{N,\gamma}_{z}(f)}.$
By Cauchy–Schwarz inequality and Young’s inequality,
$\bigl|\langle W_{\gamma,\ell,L}^{(1)},f\rangle_{\mu_{N,\gamma,\beta}}\bigr|\leq C\sum_{z\in\Lambda_{N}}|b_{z}^{\gamma,\ell,L}|\sqrt{D^{\gamma,N}_{z}(f)}\leq D_{\gamma}^{N}(f)+C\sum_{z\in\Lambda_{N}}|b_{z}^{\gamma,\ell,L}|^{2}.$ (30)
Taking the supremum over $f$ yields
$\|W_{\gamma,\ell,L}^{(1)}\|_{\mathcal{H}_{N}^{-1}}^{2}\leq C\sum_{z\in\Lambda_{N}}|b_{z}^{\gamma,\ell,L}|^{2}\leq C\|J\|_{L^{\infty}}^{2}\frac{L}{N}.$
Since $W_{\gamma,\ell,L}^{(1)}$ is centered under the reversible measure $\mu_{N,\gamma,\beta}$, the entropy–Kipnis–Varadhan estimate in Lemma 9 gives
$\mathbb{E}_{f_{N,0}}\left|\int_{0}^{T}W_{\gamma,\ell,L}^{(1)}((\sigma^{N}(\alpha_{\gamma}^{-1}t)))\,dt\right|\leq C\sqrt{NT}\,\left(C_{J}\alpha_{\gamma}\frac{L}{N}\right)^{1/2}\leq C\|J\|_{L^{\infty}}\sqrt{\alpha_{\gamma}LT}.$
Since $\alpha_{\gamma}L_{\gamma}\to 0$, we conclude that
$\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\left|\int_{0}^{T}\varepsilon_{\gamma}\sum_{i\in\Lambda_{N}}J\left(\varepsilon_{\gamma}i\right)\Bigl(m_{i}^{\ell_{\gamma}}((\sigma^{N}(\alpha_{\gamma}^{-1}t)))-\widetilde{m}_{i}^{\ell_{\gamma},L_{\gamma}}((\sigma^{N}(\alpha_{\gamma}^{-1}t)))\Bigr)dt\right|=0.$ (31)
Step 2: Sliding average versus larger block average. We now compare $\widetilde{m}_{i}^{\ell,L}$ with $m_{i}^{L}$. By direct averaging,
$\widetilde{m}_{i}^{\ell,L}(\sigma)=\frac{1}{2L+1}\sum_{|y|\leq L}\frac{1}{2\ell+1}\sum_{|u|\leq\ell}\sigma(i+y+u)=\sum_{v\in\mathbb{Z}}w_{\ell,L}(v)\sigma(i+v),$
@@
$m_{i}^{L}(\sigma)=\sum_{|v|\leq L}\frac{1}{2L+1}\sigma(i+v).$
A direct comparison of the weights shows that
$\sup_{\sigma\in\Sigma_{N}}\sup_{i\in\Lambda_{N}}\bigl|\widetilde{m}_{i}^{\ell,L}(\sigma)-m_{i}^{L}(\sigma)\bigr|\leq C\,\frac{\ell}{L}.$